OpenStudy (anonymous):

Which is the prime factorization of 98? A.2 • 7 B.2^2 • 7 C.2 • 7^2 D.2^2 • 7^2

OpenStudy (anonymous):

Break 98 into factors and then, each factor by factor to get.. 98 = 2 * 49 = 2 * 7 * 7 = 2 * 7²

OpenStudy (anonymous):

thanks try not to do all the work for me next time....I'm just warning you lol